# A 3.3904-Competitive Online Algorithm for List Update with Uniform Costs

### v1

@@ Appendix A Handling insertions and deletions
In the “dynamic” variant of List Update, insertions and deletions of items are allowed, and cost exactly as much as accessing the item. Crucially, insertions are allowed only at the end of the list; in $\textsf{LUP}_{0}$, the free exchange(s) then allow to immediately move such item to any position for free. As we demonstrate below, any algorithm can be extended to handle insertions while retaining its competitive ratio. [^3] If that ratio depends on the length of the list, i.e., equals $R(\ell)$ for a list of initial (if deletions are allowed) length $\ell$, then $\ell$ equal to the maximum length of the list at any time should be used to determine the ratio. Likewise, any algorithm can be extended to handle deletions but it appears that the effect on its competitive ratio requires further study. Thus, for deletions, we merely demonstrate that Fpm retains its competitive ratio $R$.
Insertions
For any algorithm that does not support insertions (though may support deletions), insertion of $z$ can be treated as an access operation to the item $z$ that was always at the position corresponding to the end of the list immediately after the insertion and was neither accessed nor involved in any exchange before. Specifically, if item $z$ is the $i$-th one to be inserted, then it is treated as having resided exactly $i$ positions after the last item of the initial list. We note in passing that the inserted item $z$ is treated as a new copy of $z$, particularly in the case that $z$ was in the list at any earlier time, either at the very beginning or due to being inserted (possibly multiple times) since then, and that for any prior operation on $z$, the copy which it pertains is trivially determined. Following the insertion, all the actions of the algorithm corresponding to the access operation described are triggered: both internal and those operating on the list, i.e., any exchanges, be they free or paid. are applied as stipulated by the algorithm for the access operation described. Evidently, the cost of handling an insertion this way costs exactly as much as the actually executed access operation.
The same reasoning applies to Opt, pair-based Opt, or the general work function. For pair-based Opt, as $z$ is never accessed prior to insertion, the work function $W^{yz}$ for the pair $\{y,z\}$ remains constant until then, and likewise the pair remains in mode $\alpha$. Generalizing to work functions, it is easy to see that at any time $t$ prior to the insertion of $z$, any state in which $z\prec y$ for some item $y$ in the list at time $t$ is not fundamental.
Now we comment on what this means for Fpm, and why the potential function is unaffected. Initially, $\Theta_{z}=z$ for all items, including those (yet) to be inserted. If $z$ is such an item, then for any item $y\prec z$, we have $\Theta_{y}\preceq y\prec\Theta_{z}=z$, i.e., $\{y,z\}$ pair is in state $\alpha^{d}$, contributing $\alpha^{d}=0$ to the potential. (Note that $y$ may be in the list or itself yet to be inserted.) Any item $y\succ z$ is yet to be inserted if $z$ is yet to be inserted, so we considered all pairs involving some item yet to be inserted. Moreover, since before $z$ is inserted, all the items that are accessed (incl. inserted) or possibly deleted satisfy $y\prec z$ and consequently $\Theta_{y}\prec\Theta_{z}=z$, target cleanups made before the insertion of $z$ neither alter $\Theta_{z}=z$ nor make it a target of another element. Hence, until $z$ is inserted, all the pairs involving $z$ remain in $\alpha^{d}$ state, which contributes 00 to the potential. We note that $\Theta_{z}$ is first changed (to the front of the list) once $z$ is inserted, since target cleanup, movement, and target reset are all performed as the actions corresponding to the access operation.
Deletions
To execute a deletion of $z$, any algorithm simply needs to access $z$ and remove it from the list; in particular, any exchange to be executed immediately after the access that would involve $z$ (which, for all known algorithms, means all the exchanges connected with the access) are obviously not performed. Accounting for the cost is more subtle, because a deleted item $z$ may appear at any position in the list (which may be different for the algorithm and the optimum solution): Even though the cost of deleting $z$ is no different from the cost of accessing it (ignoring aforementioned exchanges that do not take place), once $z$ is deleted, this affects the cost of all subsequent operations on any item(s) $y$ such that $y\succ z$ immediately before $z$’s deletion. Therefore, in terms of costs, this is not equivalent to simply marking $z$ as deleted and ignoring it henceforth. This is why rather than making a general statement, we only argue that Fpm remains $R$-competitive.
To clarify that Fpm is well defined, let us consider work functions first, in order to then focus on their special case, the pair-based Opt. For general work functions, the domain changes (shrinks) after a deletion, since there are fewer elements on the list to be permuted, but the work function definition (and thus the way to compute it) remain unchanged. For the work functions on pairs, involved in cost of pair-based Opt, there are clearly no changes for pairs which do not involve the deleted item. For those which do contain the deleted item $z$, the value of the function is first computed as for a regular access to $z$ and then its value remains frozen forever, since once $z$ is removed from the list, only the other item of the pair remains, never generating any cost. Moreover, we stress that while clearly there is no movement of $z$ by Fpm when it is accessed for deletion and no need to reset $\Theta_{z}$, target cleanup is still (and must be) performed: If any $y\neq z$ had $\Theta_{y}=z$ before the deletion, then after the deletion $\Theta_{y}$ points to $z$’s successor. Clearly, for these to occur, due to the $\Theta_{y}\preceq y$ invariant, $y\succ z$ has to hold, which is why the invariant remains true for $y$ after the deletion.
Hence, when Fpm is to delete $z$, we need to account for the access cost to $z$ (without exchanges) the usual way, taking the change of potential into account. The change of pair potentials is easy to track: For all pairs involving $z$, their state changes to non-existent and stays that way forever since afterwards $z$ is no longer in the list. This means that the pair potential for such pairs changes to 00 regardless of its prior value. Since the prior value of the pair potential is non-negative, the change of potential is at most 00, and as pair-based Opt values never decrease, this implies that 4.13 still holds. As observed before, the target reset may alter some $\Theta_{y}$, changing the state of some pairs that do not involve $z$. But as nothing else changes for those pairs, such changes are still characterized by 4.9 and Table 3, so 4.15 still holds for them.
Thus, to complete the proof that Fpm remains $R$-competitive, it remains to demonstrate that 4.12 continues to hold. Inspecting the proof, one can see that it suffices to show that $G_{\mathrm{del}}\odot V\leq R\cdot G_{\mathrm{OPT}}\odot V$, where $V$ and $G_{\mathrm{OPT}}=\left[\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},\textstyle{\frac{1}{2}},0\right]$ are as defined before, and $G_{\mathrm{del}}$ is the row vector which captures the cost of deleting $z$ (i.e., accessing $z$, without moving it afterward) and the corresponding change of potential. Defining $G_{\mathrm{del}}$ is easy: accessing $z$ costs $1$ per $\{y,z\}$ pair such that $y\prec z$, and for each such pair the potential changes to 00, so its change is the negative prior value, where those can be read in the first row of Table 1. Thus,
$G_{\mathrm{del}}=\left[1,1,1,1,1,1\right]-\left[0,\beta^{d},\alpha^{oe},\beta^{o},\beta^{ne},\gamma^{ne}\right]$
It is straightforward to observe that for our choice of parameters, $G_{\mathrm{del}}\odot V\leq R\cdot G_{\mathrm{OPT}}\odot V$ does indeed hold: All entries of $G_{\mathrm{del}}$ are at most $1$ because all (subtracted) pairwise potentials are non-negative, moreover $R\geq 3$ and all but the last entry of $G_{\mathrm{OPT}}$ are ${\frac{1}{2}}$. The last entry of $G_{\mathrm{OPT}}$ is 0, but the corresponding one in $G_{\mathrm{del}}$ is $11-\gamma^{ne}=-$ 1.
Appendix B Computational Results

### v3

@@ Appendix A Handling insertions and deletions
In the “dynamic” variant of List Update, except from access requests, the input $\sigma$ may contain insertions and deletions of items. By the variant definition (see, e.g., [3]):
  • inserting $z^{*}$ places it at the end of the list, and its cost is defined to be equal to the current length of the list;
  • deleting $z^{*}$ involves accessing $z^{*}$ (at a cost equal to its position), and then removing it.
Right after the insertion, an algorithm may reorganize the list paying the usual cost of swaps. Similarly, such reorganization is possible right after accessing the item to be deleted, and before its actual removal.
In this section, we demonstrate that Fpm can handle insertions and deletions while retaining its competitive ratio. We start with a few definitions and clarifications, showing how the arguments for lower-bounding Opt need to be adapted for the dynamic variant. Later, we we specify what swaps Fpm performs due to insertions and deletions.
Adjusted cost.
Let adjusted cost of an algorithm be the cost that omits the cost of serving insertions of items. (It does involve the cost of an optional list reorganization). As these costs are the same for both Fpm and Opt, if we can show that an algorithm is $R$-competitive with respect to the adjusted cost, then it is also $R$-competitive with respect to the actual cost. From this point on, we use only adjusted cost in our analysis.
Lifetimes of pairs.
Without loss of generality, we assume that when an item is deleted, it is never inserted back to the list (we simply rename its next occurrences). In the main part of the paper, we used $\mathcal{P}$ to denote the set of all pairs of items. For the dynamic variant, we will use $\mathcal{Q}$ to denote the set of all pairs of items that are ever together in the list. We call a pair $\{x,y\}\in\mathcal{Q}$
  • dormant before $x$ and $y$ are both present in the list;
  • active when both $x$ and $y$ are present in the list;
  • inactive after either $x$ or $y$ is deleted from the list.
Each pair of $\mathcal{Q}$ is active within a certain period. This period may be preceded by the dormant one, and may be followed by the inactive one. (In the static case, all pairs are always active.)
Pair-based OPT.
Previously, we related both Fpm and Opt to optimal algorithms for sequences $\sigma_{xy}$ operating on two-item lists containing $x$ and $y$. In the dynamic variant, we define $\sigma_{xy}$ to contain requests only from the period when the pair $\{x,y\}$ is active. The corresponding two-item list is also defined only within this period. The initial order of items on this two-item list matches the order of $x$ and $y$ on the actual list at the beginning of the active period. This implies that each pair starts its active period in mode $\alpha$.
With these definitions in place, 2.1 extends in a straightforward way to the dynamic case, i.e., $\sum_{\{x,y\}\in\mathcal{Q}}\textsl{{Opt}}(\sigma_{xy})\leq\textsl{{Opt}}(\sigma)$ for each input sequence $\sigma$. This also implies a counterpart of 2.2 for the dynamic variant: for any input sequence $\sigma$ consisting of $T$ requests, it holds that $\sum_{t=1}^{T}\sum_{\{x,y\}\in\mathcal{Q}}\Delta_{t}w^{xy}\leq\textsl{{Opt}}(\sigma)$.
Potential function.
Recall that the potential function $\Phi_{xy}$ for a pair $\{x,y\}$ depends on the corresponding work function $W^{xy}$, which is defined only for active pairs. Therefore, for completeness, we set the potential function for a dormant pair to be zero, and assume that the value of the potential function is frozen once a pair stops being active and becomes inactive.
Specifying the behavior of the algorithm.
The behavior of Fpm for access requests remains the same as in the static case. For insertions, Fpm performs no list reorganization. Finally, suppose that an element $z^{*}$ is deleted. Upon such request, Fpm performs exactly the same list reorganization as if $z^{*}$ was accessed (i.e., it moves $z^{*}$ fully or partially towards the front of the list). Once this action is performed, due to the target cleanup operation, $z^{*}$ is not a target for any other item in the list.
Admittedly, the behavior of Fpm for deletions is wasteful as it moves the element which is about to be removed from the list. However, this definition streamlines the analysis, as we can reuse the arguments for access requests.
Competitive ratio.
We first argue about potential changes due to insertions and deletions. Later, we use these arguments to show that Fpm retains its competitive ratio also in the dynamic variant.
We split the action of deleting an item $z^{*}$ into two parts: accessing the item (with induced list reorganization) and actual removal ot this item from the list.
Lemma A.1.
Fix a pair $\{x,y\}\in\mathcal{Q}$. Due to an insertion or an actual removal, $\Delta\Phi_{xy}=0$.
Proof A.2.
Let $z^{*}$ be the item either inserted or actually removed; we consider two cases.
In the first case, pair $\{x,y\}$ does not contain $z^{*}$. If pair $\{x,y\}$ remains dormant or inactive, then $\Phi_{xy}$ is trivially unaffected. Otherwise, the pair must be and remain active. It suffices to note that neither the insertion of $z^{*}$ nor the actual removal of $z^{*}$ does affect $\theta_{x}$ or $\theta_{y}$. Thus, the mode and flavor of $\{x,y\}$ remains unchanged, and hence $\Delta\Phi_{xy}=0$.
In the second case, one of pair elements, say $x$, is equal to $z^{*}$.
  • For insertions, this means that pair $\{z^{*},y\}$ switches from the dormant state (of zero potential) to the active one. As noted above, $\{z^{*},y\}$ starts in mode $\alpha$. Since Fpm sets $\theta_{z^{*}}=z^{*}$, it holds that $\theta_{y}\preceq y\prec\theta_{z^{*}}=z^{*}$, i.e., pair $\{z^{*},y\}$ starts in flavor $d$. Hence, the state of $\{z^{*},y\}$ right after the insertion is $\alpha^{d}$ which corresponds to the zero potential. Thus, $\Delta\Phi_{xy}=0$.
  • For actual removals, pair $\{z^{*},y\}$ switches from the active state to the inactive one. The pair potential becomes frozen, and thus $\Delta\Phi_{xy}=0$.
Theorem A.3.
The competitive ratio of Fpm in the dynamic variant is at most $R=\textstyle{\frac{1}{8}}(23+\sqrt{17})\leq 3.3904$.
Proof A.4.
Fix any sequence $\sigma$ consisting of $T$ requests. By Theorem 4.17, for any step $t$ containing an access request, we have that
$\Delta_{t}\textsl{{Fpm}}+\sum_{\{x,y\}\in\mathcal{Q}}\Delta_{t}\Phi_{xy}\leq R\cdot\sum_{\{x,y\}\in\mathcal{Q}}\Delta_{t}w^{xy}.$ (4)
We observe that the initial potentials for all pairs of $\mathcal{Q}$ (including the dormant ones) are zero. Thus, to complete the proof, it remains to show that (4) holds for insertions and deletions. The competitive ratio of Fpm then follows by summing (4) over all steps $t$, using the non-negativity of pair potentials, and finally applying the relation $\sum_{t=1}^{T}\sum_{\{x,y\}\in\mathcal{Q}}\Delta w^{xy}\leq\textsl{{Opt}}(\sigma)$ as in the proof of Theorem 4.17.
For insertions, by the definition of adjusted cost, we have that $\Delta_{t}\textsl{{Fpm}}=\Delta_{t}\textsl{{Opt}}=0$. By A.1, $\Delta_{t}\Phi_{xy}=0$ for any pair $\{x,y\}\in\mathcal{Q}$, and thus (4) holds.
For deletions, we can split the action into two parts: accessing the item (and the resulting list reorganization) and actual removal of the item. For the first part, (4) holds by Theorem 4.17. For the actual removal, $\Delta_{t}\textsl{{Fpm}}=\Delta_{t}\textsl{{Opt}}=0$. Furthermore, A.1 implies that $\Delta_{t}\Phi_{xy}=0$ for any pair $\{x,y\}\in\mathcal{Q}$, and thus (4) holds.
Appendix B Lower Bounds on MTF and its Modifications
In this section, we show the lower bounds greater than $3$ for several natural algorithms, most of which have previously been considered for this problem. All these bounds (unless explicitly stated) hold for the full cost model, and thus also for the partial one.
B.1 Specific Algorithms
Deterministic Bit.
Algorithm Deterministic Bit (Dbit) starts with all items unmarked. On a request to item $r$, if it is marked, it moves $r$ to the front and unmarks it. Otherwise, it marks $r$. We show that the competitive ratio of Dbit is at least $3.302$.
To this end, consider a list of length $n$ and denote its items as $x_{0},x_{1},\ldots,x_{n-1}$, counting from the front of the list. We divide the list into two parts, $A$ and $B$, where part $A$ has length $cn$ and part $B$ has length (1c)n(1-c)n. The value of parameter $c$ will be determined later. Formally $A=(x_{0},x_{1},x_{2},\ldots,x_{cn-1})$ and $B=(x_{cn},x_{cn+1},\ldots,x_{n-2},x_{n-1})$. For succinctness, we denote such order of the list as $A,B$. We define input
$\sigma=($ $x_{cn-1},x_{cn-2},\ldots,x_{1},x_{0},$
$x_{n-1},x_{n-1},x_{n-2},x_{n-2},x_{n-3},x_{n-3},\ldots,x_{cn},x_{cn})^{2}.$
In this sequence, all items of $A$ are requested first, starting from the one furthest in the list. Then every item of $B$ is requested twice, also starting from the furthest one. It results in items of $B$ being transported to the front, thus after the first half of $\sigma$, the order of the list of Dbit is $B,A$. Note that order of items within $B$ does not change.
Then the same sequence of requests is repeated. The difference is that now requesting any item of $A$ results in moving that item to the front. Then items of $B$ are again requested twice, resulting in final list of algorithm being $B,A$. Then,
$\textsl{{Dbit}}(\sigma)$ $=\frac{cn(cn+1)}{2}$ (access to $A$)
$\quad+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$\quad+cn\cdot(2n-1)$ (access to $A$ and movement)
$\quad+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$=n^{2}\left(6-4c+\frac{c^{2}}{2}\right)+n\left(\frac{3c}{2}-2\right).$
Now we define a better algorithm for this sequence, which upper-bounds the cost of Opt. After receiving the first batch of requests to the items of $A$, it moves all the items of $B$ before the items of $A$. It does not performs any further swaps. Thus,
$\textsl{{Opt}}(\sigma)$ $\leq\frac{cn(cn+1)}{2}$ (access to $A$)
$\quad+cn\cdot(1-c)n$ (swap $A$ with $B$)
$\quad+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$\quad+\frac{(1-c)n+1+n}{2}\cdot cn$ (access to $A$)
$\quad+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$=n^{2}\left(2-2c+c^{2}\right)+n\left(2-c\right).$
Note that the final order of the list in both algorithms is the same and equal to $B,A$. Moreover, every item is requested an even number of times, so no item is marked after $\sigma$. Thus, this request sequence can be repeated infinitely many times, with sets $A,B$ redefined after each sequence.
With a long enough list, $n$ becomes insignificant compared to $n^{2}$, and therefore the ratio is $\textsl{{Dbit}}(\sigma)/\textsl{{Opt}}(\sigma)\geq(6-4c+\frac{1}{2}c^{2})/(2-2c+c^{2})$. For $c=\frac{1}{3}(5-\sqrt{13})$, we have $\textsl{{Dbit}}(\sigma)/\textsl{{Opt}}(\sigma)\geq\frac{1}{2}(3+\sqrt{13})\geq 3.302$.
Deterministic Bit in partial cost model.
We now show that the competitive ratio of Dbit is at least $4$ in the partial cost model. We recursively construct a sequence that achieves this ratio for every length of the list $n\geq 2$. For a list of length $n$, the created sequence will be denoted $\sigma_{n}$. For reasons that will become apparent soon, we index items from the end of the list, namely the algorithm’s list at the start of the sequence is $x_{n-1},x_{n-2},\ldots,x_{1},x_{0}$.
First, assume $n=2$. Set
$\sigma_{2}=x_{1},x_{0},x_{0},x_{1},x_{0},x_{0}.$
Clearly, $\textsl{{Dbit}}(\sigma_{2})=8$. The cost of $2$ can be achieved for that sequence by starting with the list order $x_{1},x_{0}$ and swapping items only after the first request. Note that such an algorithm would have the same order of items as algorithm Dbit after serving $\sigma_{2}$.
Now let $n>2$. Set
$\sigma_{n}\;=\;x_{n-1},\sigma_{n-1},x_{n-1},\sigma^{\prime}_{n-1}.$
where by $\sigma^{\prime}_{n-1}$ above we mean the sequence obtained from $\sigma_{n-1}$ by permuting the items $x_{n-2},\dots,x_{0}$ according to their position in the list after executing $\sigma_{n-1}$. Specifically, if $x_{\pi(n-2)},x_{\pi(n-3)},\dots,x_{\pi(0)}$ is the list after executing $\sigma_{n-1}$ then $\sigma^{\prime}_{n-1}$ is obtained from $\sigma_{n-1}$ by replacing each item $x_{i}$ by $x_{\pi(i)}$. For example, for $n=3$,
We now express $\textsl{{Dbit}}(\sigma_{n})$ and $\textsl{{Opt}}(\sigma_{n})$ as functions of $\textsl{{Dbit}}(\sigma_{n-1})$ and $\textsl{{Opt}}(\sigma_{n-1})$, respectively.
At the start of $\sigma_{n}$, item $x_{n-1}$ is at the front, marked because of the first request. Within $\sigma_{n-1}$ every item other than $x_{n-1}$ is requested even number of times (at least twice). For the first pair of such requests, $x_{n-1}$ contributes $3$ to the cost (one per each request and one for the swap when the requested item is moved to the front). Thus, the additional cost due to $x_{n-1}$ is at least 3(n1)3(n-1). Then, item $x_{n-1}$ is requested and moved to the front incurring cost 2(n1)2(n-1). Similarly, $x_{n-1}$ contributes an additional cost of 3(n1)3(n-1) to $\sigma^{\prime}_{n-1}$. Thus,
The optimal algorithm can serve the first request to $x_{n-1}$ at cost 0 and then move it immediately to the end, paying 2(n1)2(n-1) for the second request to $x_{n-1}$. In the final list of Opt, $x_{n-1}$ is at the end, just like in the list of Dbit. This implies that serving $\sigma_{n-1}$ and $\sigma^{\prime}_{n-1}$ will cost at most $2\cdot\textsl{{Opt}}(\sigma_{n-1})$, i.e.,
$\textsl{{Opt}}(\sigma_{n})\;\leq\;2\cdot\textsl{{Opt}}(\sigma_{n-1})+2\cdot(n-1).$ (6)
Half-Move.
Finally, we consider the algorithm Half-Move, which moves the requested item to the middle point between its position and the front of the list. We show that its competitive ratio is at least $6$.
Consider a list of even length $n$. Denote the items as $x_{0},x_{1},\ldots,x_{n-1}$, according to their positions in the initial list of the algorithm. The request sequence is defined as
$\sigma=\left(x_{n-1},x_{n-2},x_{n-3},\dots,x_{n/2}\right)^{k}$
for some $k\geq 1$. At the time of the request, the requested item is at the end of the list and it is moved to the middle of the list. Since in the list of even length there are two middle positions, we will assume that the item is always moved to the one further from the front of the list. (If it is moved to the other one, the proof only needs to be slightly modified.) The cost of each request for the algorithm is $n+\frac{1}{2}n-1=\frac{3}{2}n-1$. Moreover, after $\frac{1}{2}n$ requests, the list returns to its original order. Therefore, the total cost of the sequence is $k\cdot\frac{1}{2}n\cdot(\frac{3}{2}n-1)$.
It is possible to change the order of the list to $x_{n/2},x_{n/2+1},\ldots,x_{n-1},x_{0},x_{1},x_{2},\ldots,x_{n/2-1}$ before any requests arrive, at cost less than $n^{2}$. Thus, the total cost of the request sequence is $k\cdot\binom{n/2}{2}$. Thus, $\textsl{{Opt}}(\sigma)\leq n^{2}+\frac{1}{8}kn^{2}$. The ratio approaches $6$ for sufficiently large $n$, as $k\rightarrow\infty$.
B.2 Stay-or-MTF Class
We now consider a class of online algorithms that never make partial moves. That is, when an item $x$ is requested, an algorithm either does not change the position of $x$ or moves $x$ to the front. We refer to such algorithms as Stay-or-MTF algorithms.
We show that if $A$ is a Stay-or-MTF algorithm, then the competitive ratio of $A$ is at least $3.25$, even for a list of length $3$.
The adversary strategy is illustrated in the figure below. The list has three items named $a$, $b$ and $c$. The vertices represent states of the game. In each state, we assume that the $A$’s list is abcabc. When $A$ executes a move, the items are appropriately renamed. Each state is specified by the current offset function, equal to the current work function minus the minimum of this work function. At each move, the increase of the minimum of the work function is charged towards the Opt’s cost, and the offset function is updated. Each offset function that appears in this game is specified by the set of permutations of $a,b,c$ where its value is 0. (Its value at each other list is equal to the minimum swap distance to some offset function with value 0.) To specify this set, we indicate, using braces, which items are allowed to be swapped. For example, $a{\left\{{bc}\right\}}$ represents two permutations, abcabc and acbacb, and $\left\{{abc}\right\}$} represents all permutations.
The adversary strategy is represented by the transitions between the states. For each state the adversary specifies the generated request when this state is reached. Notation $r/\xi$ on an arrow means that the request is $x$ and the adversary cost is $\xi$. Then, the algorithm decides to make a move. The decision points of the algorithm are indicated by small circles. The type of move is uniquely specified by the cost, specified on the edges from these circles. For example, if the request is on $c$, the algorithm either stays, which costs $2$, or moves to front, which costs $4$.
The numbers next to each state are the values of potential $\Psi$. By routine verification of the transition graph, we can check that for each move we have
$\Delta A+\Delta\Psi\;\geq\;\textstyle{\frac{13}{4}}\cdot\Delta\textsl{{Opt}},$
where $\Delta A$, $\Delta\Psi$ and $\Delta\textsl{{Opt}}$ represent the cost of $A$, the potential change, and the optimum cost for this move. This shows that the competitive ratio of $A$ is at least ${\frac{13}{4}}=3.25$.
Our construction is tight for $n=3$, i.e., it can be shown that there exists a Stay-or-MTF algorithm with the competitive ratio of $3.25$ for $n=3$. For $n=4$, we can improve the lower bound to $3.3$, using an approach that is essentially the same, but requires many more states.
Appendix C A 3-Competitive Algorithm for Lists of Length 3
The algorithm Fpm is $3$-competitive for lists of length 3. This is true even against our lower bound $\sum_{\{x,y\}\in\mathcal{P}}w^{xy}$ on the optimum cost.
To prove that, we consider possible states of Fpm. There are 15 of them. We assign a potential $\Psi$ to every state such that for any state and for any request, it holds that $\Delta\textsl{{Fpm}}+\Delta\Psi\leq 3\cdot\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$. Note that $\Psi$ is introduced only to simplify the proof: it is not related to the potential used internally by the algorithm Fpm.
The states are shown in Table 4. In each state we assume that the list is abcabc; the items are renamed after a transition if needed. Arrows represent modes: an arrow pointing from an item $x$ towards an item $y$ indicates that the pair $\{x,y\}$ is in mode $\alpha$ (if $x$ is before $y$ in the list) or $\gamma$ (otherwise). If there is no arrow between $x$ and $y$, the pair is in mode $\beta$. The column “move” specifies what Fpm does after the request, while the column “dest.” denotes the index of the destination state after movement. State changes correspond to the moves of Fpm. Cost of the pair-based Opt is denoted $\Delta\textsl{{Opt}}_{P}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$.
state idx $\Psi$ req. move dest. $\Delta\textsl{{Fpm}}$ $\Delta\Psi$ $\Delta\textsl{{Opt}}_{P}$
[Uncaptioned image] 0 0 $a$ Front 0 0 0 0
$b$ $\Theta_{b}$ 11 11 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ 22 22 11 11
[Uncaptioned image] 11 ${\frac{1}{2}}$ $a$ Front 0 0 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 0 22 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ 33 22 11 11
[Uncaptioned image] 22 11 $a$ Front 44 0 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ 55 11 11 11
$c$ Front 0 44 1-1 11
[Uncaptioned image] 33 $\frac{3}{2}$ $a$ Front 44 0 1-1 11
$b$ Front 44 22 1-1 11
$c$ Front 44 44 1-1 11
[Uncaptioned image] 44 ${\frac{1}{2}}$ $a$ Front 44 0 0 0
$b$ $\Theta_{b}$ 11 11 0 11
$c$ $\Theta_{c}$ 11 33 0 11
[Uncaptioned image] 55 22 $a$ Front 66 0 0 11
$b$ Front 44 22 $\frac{3}{2}$ ${\frac{1}{2}}$
$c$ Front 77 44 1-1 11
[Uncaptioned image] 66 22 $a$ Front 66 0 0 0
$b$ $\Theta_{b}$ 11 11 $\frac{3}{2}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ 88 33 0 11
[Uncaptioned image] 77 11 $a$ Front 44 0 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ 99 11 11 11
$c$ Front 0 44 1-1 11
[Uncaptioned image] 88 22 $a$ Front $10$ 0 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 0 22 2-2 11
$c$ $\Theta_{c}$ $11$ 22 11 11
[Uncaptioned image] 99 22 $a$ Front 0 0 2-2 11
$b$ Front $10$ 22 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $12$ 22 11 11
[Uncaptioned image] $10$ $\frac{3}{2}$ $a$ Front $10$ 0 0 0
$b$ $\Theta_{b}$ 11 11 1-1 11
$c$ $\Theta_{c}$ $13$ 22 11 11
[Uncaptioned image] $11$ 33 $a$ Front $14$ 0 1-1 11
$b$ Front 77 22 2-2 11
$c$ Front 44 44 $\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $12$ 33 $a$ Front 77 0 2-2 11
$b$ Front $14$ 22 1-1 11
$c$ Front 44 44 $\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $13$ $\frac{5}{2}$ $a$ Front $14$ 0 ${\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ 33 11 1-1 11
$c$ Front 0 44 $\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $14$ 22 $a$ Front $14$ 0 0 0
$b$ $\Theta_{b}$ 88 11 0 11
$c$ $\Theta_{c}$ 11 33 $\frac{3}{2}$ ${\frac{1}{2}}$
Table 4: Fpm is 3-competitive for lists of length 3
There exists an easier $3$-competitive algorithm for $3$ items. This algorithm uses $6$ states to keep track of the past history. The algorithm is illustrated below as a graph whose vertices represent the current state of the algorithm and edges represent its transitions. Again, items are renamed after each transition so that the list is always abcabc. The modes of each pair are represented the same way as in Table 4. Each edge is labeled by the requested item, the algorithm’s cost, and the optimum cost (${\frac{1}{2}}$ for each mode change). The algorithm’s move is implied by the cost value. For example, if the request is $c$ and the cost is $3$, it means that the algorithm paid $2$ to access $c$ and then swapped $c$ with the preceding $b$.
Appendix D Lower bound for the Full-Or-Partial-Move algorithm
In this section, we show that for the choice of parameters from 4.12, the competitive ratio of Fpm is at least $3.04$ in the partial cost model.
The sequence achieving this ratio is shown in Table 5. It uses five items (on a longer list one may simply use only five initial items). Graphs in the column “work function on pairs” depict modes of all the pairs. There, an arrow from $x$ to $y$ means that $W^{xy}(xy)+1=W^{xy}(yx)$ (which corresponds to mode $\alpha$ or $\gamma$, depending on the order of the list). Some arrows are omitted for clarity: the actual graph is the transitive closure of the graph shown.
Table 5 compares cost of Fpm to the cost of pair-based Opt defined as $P_{OPT}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$, which for this sequence is equal to $25$. We now show the behavior of an actual offline algorithm Opt that achieves this cost. Before the start of the sequence, the algorithm changes its list to cdeabcdeab. This incurs an additional additive cost, which becomes negligible if the sequence is repeated sufficiently many times. During the entire sequence, Opt performs two moves: after serving the first request, it moves $c$ to the third position in the list, and after the ninth request, it brings $c$ back to the front. Thus, it executes $4$ swaps. It is easy to check that the cost of requests for paid by Opt is $21$, and hence its total cost is $25$.
list of Fpm work function req. move cost cost
on pairs of Fpm of Fpm of $\textsl{{Opt}}_{P}$
[Uncaptioned image] [Uncaptioned image]
$c$ $\theta_{c}$ 2 $1\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$e$ $\theta_{e}$ 6 $3\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$e$ Front 14 $5\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$d$ $\theta_{d}$ 20 $7\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$c$ Front 28 $9\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$d$ $\theta_{d}$ 33 11
[Uncaptioned image] [Uncaptioned image]
$e$ $\theta_{e}$ 35 12
[Uncaptioned image] [Uncaptioned image]
$e$ Front 39 $12\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$d$ $\theta_{d}$ 42 $13\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$c$ $\theta_{c}$ 44 $14\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$c$ Front 48 $15\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$b$ $\theta_{b}$ 52 $17\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$b$ Front 60 $19\textstyle{\frac{1}{2}}$
[Uncaptioned image] [Uncaptioned image]
$d$ $\theta_{d}$ 64 21
[Uncaptioned image] [Uncaptioned image]
$a$ $\theta_{a}$ 68 23
[Uncaptioned image] [Uncaptioned image]
$a$ Front 76 25
[Uncaptioned image] [Uncaptioned image]
Appendix E Computational Results